OpenStudy (anonymous):

which expression is equivalent to 4x√3:(81) a.((4)/(3))x9^3 b. 4x(9)^(1/3) c.((4)/(3))x81 d.((4)/(3))x(9)^(1/2) e. 4x(9)^(2/3) The anwser is E but i dont know how to get it

OpenStudy (rane):

r u sure there is a square root in the q?

OpenStudy (anonymous):

what q

OpenStudy (anonymous):

its a 3

OpenStudy (rane):

4x√3:(81) r u sure there is a square root

OpenStudy (anonymous):

its 4 X 3 square root 81

OpenStudy (rane):

the answer is E they both give u 108

OpenStudy (rane):

\[4 X 3 \sqrt{81} = 108\]

OpenStudy (rane):

also 4 X 9 ^2/3 = 108

OpenStudy (anonymous):

well i know that 4 X 9^2/3 = 108 but i dont get how you know thats the anwser

OpenStudy (rane):

use calculator

OpenStudy (anonymous):

ty

OpenStudy (rane):

yw:)